OpenStudy (anonymous):

Please Solve by usingsubstitution: y=5-x 1=4x+3y

OpenStudy (anonymous):

X=14 Y=-9

OpenStudy (anonymous):

can you please tell me how you got that?

OpenStudy (anonymous):

yes no problem

OpenStudy (anonymous):

first you have to have y by its self in both expressions. Y=5-X IS FINE but 1=4X + 3Y has to be changed to -3Y=4X-1 THEN PLUG ONE IN TO THE OTHER WHICH NOW ITS -3(5-X)=4X-1 -15+3X=4X-1 NEXT COMBINE LIKE TERMS -15+1=4X-3X -14=X nOW ONCE YOU HAVE THAT PLUG IN TO THE OTHER EXPRESSION Y=5-(-14) Y=19 SO YOUR ANSWERS ARE X=-14 AND Y=19 SORRY WORTE IT WORNG THE FIRST TIME I GAVE YOU THE ANSWER

OpenStudy (anonymous):

ok, thanks!

OpenStudy (anonymous):

ur welcome